OpenStudy (anonymous):

What is the slope of the line 2x + 5y = -10? Choose one answer. a. negative five halves b. five halves c. two fifths d. negative two fifths e. none of the above

OpenStudy (anonymous):

Isolate for y, then put the equation into the slope-intercept form (y=mx+b). The value of m is your slope.

OpenStudy (the_fizicx99):

2x + 5y = -10 -2x -2x 5y = -2x - 10 divide every term by 5 to isolate y 5y = -2x -10 /5 /5 /5 y = -2/5x - 2 Now you have y = mx + b The slope is M The y intercept is B

OpenStudy (the_fizicx99):

Can you figure it out?

OpenStudy (anonymous):

HAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHA- no. :/

OpenStudy (the_fizicx99):

._. In y = mx + b The slope is M Then y intercept is B So, you have y = -2/5x -2 What is the slope?

OpenStudy (anonymous):

i have no idea lol

OpenStudy (anonymous):

c?

OpenStudy (the_fizicx99):

-2/5 is ?

OpenStudy (anonymous):

idk im just guessing! :(

OpenStudy (the_fizicx99):

Please try.